OpenStudy (dinorap1):

4 math questions, please check my answers. Will give medal to best response.

Arcadiouse (arcadiouse):

SUre

Arcadiouse (arcadiouse):

Well, I do not know about question 1 as we do not know what the lesson was about. Question 2: Correct Questin 3: IDK Question 4: idk

Arcadiouse (arcadiouse):

I think question 3 is correct too.

OpenStudy (dinorap1):

For number 1, you don't really need to have gone through the lesson to know the best answer. For number 1, I think the answer can either be B or D.

Arcadiouse (arcadiouse):

Sorry that I couldnt help you with question 4. Oh, thats great.

OpenStudy (dinorap1):

What did you do for number 3? Since it's only asking for overages, it's most likely D. I did: Step 1: 0.35*40=14 Step 2: 14*1.065=14.91

Arcadiouse (arcadiouse):

Real quick, where did you get the 1.065?

OpenStudy (dinorap1):

That's the percentage: 6.5%

Arcadiouse (arcadiouse):

No one needed.... Its 6.5% so .065..

Arcadiouse (arcadiouse):

Unless you can explain why you added the one?

OpenStudy (dinorap1):

Another way to get 14.91: Step 1: 0.35*40=14 Step 2: 0.065*14=0.91 Step 3: 0.91+14=14.91 I added the 1 because it makes it simpler, and gets rid of the extra step.

Arcadiouse (arcadiouse):

Yeah. You're right. The second why is how I got mine.

OpenStudy (dinorap1):

Okay, so number 3 is probably right then. The only way I could get this wrong is if they wanted the overages+everything else

OpenStudy (dinorap1):

4 is even easier, I think. Here's how to do it: Step 1: 25.95*12=311.40 Step 2: 311.40*1.045=325.41 Sorry for taking so long, I kept getting distracted irl

OpenStudy (dinorap1):

@Arcadiouse

Arcadiouse (arcadiouse):

Oh it is good. I got nothing to do anyway. thought I should help some people on here. No problem! Thanks for the metal too.

OpenStudy (dinorap1):

You're welcome! You helped me out a lot, and gave detailed explanations. That's rare on this website, so you deserve it :)

OpenStudy (dinorap1):

Want me to screenshot the 2 pages of the lesson so you can read and maybe understand #1 better?

Arcadiouse (arcadiouse):

Oh <3 I appreciate it! Sure. I would love too.

OpenStudy (dinorap1):

Page 2/5 (Page 1 is just an intro)

OpenStudy (dinorap1):

Well actually that's the only page with information on it. The rest are math and then the quick check. @Arcadiouse

Arcadiouse (arcadiouse):

Yeah, I would say the answer is right...

OpenStudy (dinorap1):

Okay! So I will probably get a 4/4. Thanks for all your help. :)

OpenStudy (dinorap1):

@Arcadiouse :)

Arcadiouse (arcadiouse):

Excellent! Good job man!

Arcadiouse (arcadiouse):

Metal for your excellent grade

OpenStudy (dinorap1):

Hahaha thanks! You earned yours too! :D Will tag you in future math quick checks definitely! Awesome help. :)

Arcadiouse (arcadiouse):

Sure. No problem with that. Thanks. Your a great "student"